The length of a rectangle is 4 m. more than the width. The area is 30 m^2. Find the width and length.
I know that the equation is w(w+4)=30, I figured out that much. I also know that you have to distribute: w^2+4w=30. But I'm so lost on what to do next.

Answer by rothauserc(4718)

We are given the following problem
The length of a rectangle is 4 m. more than the width. The area is 30 m^2. Find the width and length.
We have the following formula to solve
w(w+4) = 30
w^2+4w-30 = 0
now use quadratic formula
x = (-4 + or - sqrt(16-4(-30))) / 2
x = (-4 + or - sqrt(136)) / 2 = (-4 + or - 11.66) / 2
x = (-4 + 11.66) / 2 = 3.83
x = (-4 - 11.66) / 2 = -7.83